OpenStudy (anonymous):

For the real-valued functions f(x)=squareroot of 3x+21 and g(x)=x-2 , find the composition f(g)and specify its domain using interval notation. (f(g))(x)=? domain of f(g)=?

OpenStudy (anonymous):

\[f(x)=\sqrt{3x+21}\] \[g(x)=x-2\] \[f(g(x))=f(x-2)=\sqrt{3(x-2)+21}\]

OpenStudy (anonymous):

which of course simplifies to \[\sqrt{3x+15}\]

OpenStudy (anonymous):

domain \[3x+15\geq0\] \[3x\geq-15\] \[x\geq-3\]

OpenStudy (anonymous):

w8, its f times g (x)

OpenStudy (anonymous):

well then it would be different but you wrote \[f(g(x))\] which is a composition, not a multiplication

OpenStudy (anonymous):

(f times g) (x)

OpenStudy (anonymous):

\[fg\] is not the same as \[f(g)\]

OpenStudy (anonymous):

you wrote \[f(g(x))\] twice so i will be willing to bet that that is what it is

OpenStudy (anonymous):

btw i made am mistake. \[3x\geq-15 \] \[x\geq-5\]